OpenStudy (anonymous):

sqrt{25/64} sqrt{25/64} @Mathematics

OpenStudy (anonymous):

\[\sqrt{\left(\begin{matrix}25 \\64 \end{matrix}\right)}\]

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{\frac{25}{64}}\] \[\Large \frac{\sqrt{25}}{\sqrt{64}}\] \[\Large \frac{5}{8}\] So \[\Large \sqrt{\frac{25}{64}}=\frac{5}{8}\]